Beam OFF. Run started 20/2/22 18:42
Fast comparator threshold HEC set to 0xff for all fees and asics to combat high energy noise bursts.
Electron cooler scalar plugged back into aida03 ~3000 per sec.
System checck OK, except white rabbit all fail
Temperatures OK
Leakage current 3.638 uA.
Statistics rates in noraml range, aida03 slightly higher as expected with scalar. |